# Diophantine Approximation with Piatetski-Shapiro Primes

### v2

@@ Abstract.
We prove that for every irrational number $\alpha$, real number $\beta$, real number $c$ satisfying $81<c<9/$ 8 and positive real number $\theta$ satisfying $\theta<(9/c-8)/10$, there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
@@ 1. Introduction
In this article, we consider Diophantine approximation with denominators which are restricted to Piatetski-Shapiro primes. These are prime numbers $p$ of the form $\left[n^{c}\right]$, where $c>1$ is a fixed real number, $n$ runs over the positive integers, and $[x]$ denotes the integral part of $x\in\mathbb{R}$. One may conjecture that, given any non-integer $c>1$, there exist infinitely many such primes. Their investigation was initiated by Piatetski-Shapiro [9] who demonstrated their infinitude for $1.09091<c<12/11=$ 1.0909 …. This range has been widened by many authors. The latest record is due to Rivat and Wu [10] who obtained a range of $1.18531<c<243/205=$ 1.1853 …. Progress on Piatetski-Shapiro primes measures progress on exponential sums and sieve methods.
The Dirichlet approximation theorem, a cornerstone in Diophantine approximation, implies that for any irrational number $\alpha$, there exist infinitely many positive integers $q$ such that $||\alpha q||<q^{-1}$, where $||x||$ is the distance of $x\in\mathbb{R}$ to the nearest integer. Interesting problems arise when the $q$’s are restricted to a sparse subset of the natural numbers. Many authors have considered the question for which $\theta>0$ one can show the infinitude of primes $p$ such that $||\alpha p||<p^{-\theta}$. It may be expected that an exponent of $\theta=1-\varepsilon$ is admissible. The first author to consider this problem was Vinogradov who proved that $\theta=1/5-\varepsilon$ is admissible [11]. The current record is due to Matomäki [8] who obtained an exponent of $\theta=1/3-\varepsilon$. Harman’s sieve has become a standard tool for the investigation of this problem (for details on Harman’s sieve see [5]).
@@
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $111<c<12/$ 11 and a positive real number $\theta$ satisfying $\theta<(12/c-11)/26$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
@@
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $81<c<9/$ 8 and a positive real number $\theta$ satisfying $\theta<(9/c-8)/10$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
@@
Acknowledgements. The first-named author would like to thank the Ramakrishna Mission Vivekananda Educational and Research Insititute for an excellent work environment. The research of the second-named author was supported by a Prime Minister Research Fellowship (PMRF ID- 0501972), funded by the Ministry of Education, Govt. of India.
2. Simplification and application of Harman’s sieve
Let us denote by $\mathbb{P}$ the set of prime numbers. Our goal is to prove that
$\sharp\left\{p\in\mathbb{P}:X/2\leq p<X,\ p=\left[n^{c}\right]\mbox{ for some }n\in\mathbb{N},\ ||\alpha p+\beta||<\Delta\right\}>0$
for infinitely many $X\in\mathbb{N}$, where we set
@@ 2. Simplification and application of Harman’s sieve
Let $\gamma=1/c$. Obviously, $p=\left[n^{c}\right]$ for some $n\in\mathbb{N}$ if and only if
Moreover, $(p+1)^{\gamma}=p^{\gamma}+\gamma p^{\gamma-1}+O\left(p^{\gamma-2}\right)$. Hence, if
$\delta:=\frac{\gamma X^{\gamma-1}}{10},$ (2.3)
then
$||p^{\gamma}+2\delta||<\delta.$ (2.5)
Therefore, it suffices to show that for infinitely many $X\in\mathbb{N}$, we have
@@
where we set
$\mathcal{A}:=\left\{a\in\mathbb{N}:X/2\leq a<X,\ ||\alpha a+\beta||<\Delta,\ ||a^{\gamma}+2\delta||<\delta\right\}.$ (2.7)
By the Dirichlet approximation theorem, mentioned in section 1, there exist infinitely many positive integers $q$ such that
We shall prove that (2.6) holds if $X$ is a positive integer such that
$X^{2\theta+10\eta}\leq q\leq X^{1-\theta-10\eta}$ (2.9)
for some $q\in\mathbb{N}$ satisfying (2.8), where $\eta$ is a suitably small positive number. Since there are infinitely many such positive integers $X$, the result in Theorem 1.2 then follows.
To establish (2.6), we shall employ Harman’s sieve. Set
@@
Following usual terminology, we call the sums in (2.12) type I and the sums in (2.13) type II sums.
@@ 3. Fourier analysis
We shall detect the conditions $||\alpha a+\beta||<\Delta$ and $||a^{\gamma}+2\delta||<\delta$ in (2.7) using the following Fourier analytic device.
@@
$\chi_{\xi}(x)=\begin{cases}1,\ \text{ if }||x||<\xi,\\ 0,\ \text{otherwise.}\end{cases}$
@@
Here and in the following, the notation $x\sim X$ means that $X/2\leq x<X$. In view of Proposition 2.1 and the prime number theorem, to establish Theorem 1.2, it now suffices to prove that
@@ 4. Tailoring the type I and II sums
for $\eta>0$ small enough. Here and throughout the sequel, for $x,y>0$, the notation $x\asymp y$ means that there are constants $C_{2}>C_{1}>0$ such that $C_{1}x\leq y\leq C_{2}x$.
@@ 5. Estimations of $\mathcal{S}_{1}^{\ast}$ and $\mathcal{T}_{1}^{\ast}$
Recall (2.8) and (2.9). To estimate $\mathcal{S}_{1}^{\ast}$ and $\mathcal{T}_{1}^{\ast}$, defined in (4.1) and (4.4), we use the following standard bounds for sums involving linear exponential terms.
@@
$\sum\limits_{n\asymp N}e(\alpha kn)\ll\min\left\{N,||\alpha k||^{-1}\right\}\mbox{ for all }k\in\mathbb{N},$ (5.1)
$\sum\limits_{k\asymp K}\min\left\{N,||\alpha k||^{-1}\right\}\ll\left(\frac{KN}{q}+K+q\right)(\log 2KNq)$ (5.2)
@@
$\mathcal{S}_{1}^{\ast}\ll\sum_{m\sim M}\sum_{l\sim U}a_{m}^{\ast}c_{l}^{\ast}\max\left\{\frac{X}{M},\frac{1}{||\alpha ml||}\right\}.$
@@ 6. Estimations of $\mathcal{S}_{2}^{\ast}$ and $\mathcal{T}_{2}^{\ast}$
We first estimate the type II sum $\mathcal{T}_{2}^{\ast}$.
@@
Estimating $\mathcal{T}_{2}^{\ast}$ in a similar way as the term $L$ in [7, section 4], we have
@@
where the first line is found in [7, page 257] (with $X$ replaced by $N$, and $N$ replaced by $Y$), and for the second line we have used that $MN\asymp X$. Thus $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$ if
@@
Recalling $X^{7/22}\leq M\leq X^{8/22}$ from (4.8) and $\gamma=1/c>8/9$ from Theorem 1.2, and noting that $41-\gamma<7/22<8/22<5\gamma-$ 4 if $\gamma>8/9$, the desired bound $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$ now follows upon taking $\varepsilon$ and $\eta$ sufficiently small. ∎
@@
Lemma 6.3.
Suppose that $\varepsilon$ and $\eta$ are small enough and (6.2) holds. Then $\mathcal{S}_{2}^{\ast}\ll X^{1-2\eta}$.
@@
provided that $M\leq X^{2/3}$. This implies the result in Lemma 6.4. ∎
@@
where we have used $V\leq H\ll X^{1-\gamma+\eta}$. This implies the result in Lemma 6.5. ∎
@@
if $\gamma>8/9$, Lemmas 6.3, 6.4 and 6.5 cover all relevant ranges if $\varepsilon$ and $\eta$ are small enough, and we thus have the following.
@@
Here we point out that the condition $\gamma>8/9$ above comes from the inequality $35-5\gamma<4\gamma-$ 3.
@@ 7. Estimations of $\mathcal{S}_{3}^{\ast}$ and $\mathcal{T}_{3}^{\ast}$
Lemma 7.2.
Suppose that $\varepsilon$ and $\eta$ are small enough and (7.5) holds. Then $\mathcal{S}_{3}^{\ast}\ll X^{1-2\eta}$.

### v3

@@ Abstract.
We prove that for every irrational number $\alpha$, real number $\beta$, real number $c$ satisfying $1<c<9/8$ and positive real number $\theta$ satisfying $\theta<(9/c-8)/10$, there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $\alpha p+\beta||<p^{-\theta}$.
@@ 1. Introduction
In this article, we consider Diophantine approximation with denominators which are restricted to Piatetski-Shapiro primes. These are prime numbers $p$ of the form $\left[n^{c}\right]$, where $c>1$ is a fixed real number, $n$ runs over the positive integers, and $[x]$ denotes the integral part of $x\in\mathbb{R}$. One may conjecture that, given any non-integer $c>1$, there exist infinitely many such primes. Their investigation was initiated by Piatetski-Shapiro [9] who demonstrated their infinitude for $1<c<12/11=1.0909$.... This range has been widened by many authors. The latest record is due to Rivat and Wu [10] who obtained a range of $1<c<243/205=1.1853$.... Progress on Piatetski-Shapiro primes measures progress on exponential sums and sieve methods.
The Dirichlet approximation theorem, a cornerstone in Diophantine approximation, implies that for any irrational number $\alpha$, there exist infinitely many positive integers $q$ such that $\alpha q||<q^{-1}$, where $x||$ is the distance of $x\in\mathbb{R}$ to the nearest integer. Interesting problems arise when the $q$’s are restricted to a sparse subset of the natural numbers. Many authors have considered the question for which $\theta>0$ one can show the infinitude of primes $p$ such that $\alpha p||<p^{-\theta}$. It may be expected that an exponent of $\theta=1-\varepsilon$ is admissible. The first author to consider this problem was Vinogradov who proved that $\theta=1/5-\varepsilon$ is admissible [11]. The current record is due to Matomäki [8] who obtained an exponent of $\theta=1/3-\varepsilon$. Harman’s sieve has become a standard tool for the investigation of this problem (for details on Harman’s sieve see [5]).
@@
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $1<c<12/11$ and a positive real number $\theta$ satisfying $\theta<(12/c-11)/26$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $\alpha p+\beta||<p^{-\theta}$.
@@
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $1<c<9/8$ and a positive real number $\theta$ satisfying $\theta<(9/c-8)/10$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $\alpha p+\beta||<p^{-\theta}$.
@@
1.1. Notation
In this article, we will use the following notations.
  • We will denote the set of integers, positive integers and primes by $\mathbb{Z}$, $\mathbb{N}$ and $\mathbb{P}$, resprectively.
  • Expressions of the form f(x)=O(g(x)){f}(x)=O({g}(x)), ${f}(x)\ll{g}(x)$, and ${g}(x)\gg{f}(x)$ signify that $|{f}(x)|\leq C|{g}(x)|$ for all sufficiently large $x$, where $C>0$ is an absolute constant. A subscript of the form $\ll_{A}$ means that the implied constant may depend on the parameter $A$.
  • For $x,y>0$, the notation $x\asymp y$ means that there are constants $C_{2}>C_{1}>0$ such that $C_{1}x\leq y\leq C_{2}x$.
  • The notation $x\sim X$ means that $X/2\leq x<X$.
  • For any real number $z$, we will write $e(z):=e^{2\pi iz}$.
  • We will denote the divisor function by $\tau$, that is, for any $n\in\mathbb{N}$,
    $\tau(n):=\sum_{d|n}1.$
Acknowledgments. The authors thank the anonymous referee for his valuable comments. The first-named author thanks the Ramakrishna Mission Vivekananda Educational and Research Institute for an excellent work environment. The research of the second-named author was supported by a Prime Minister Research Fellowship (PMRF ID- 0501972), funded by the Ministry of Education, Govt. of India.
2. Application of Harman’s sieve
For $X\in\mathbb{N}$ define
[…]
where we set
@@ 2. Application of Harman’s sieve
and
$\delta:=\frac{\gamma X^{\gamma-1}}{10}.$ (2.3)
The bulk of our proof of Theorem 1.2 will consist of establishing the following.
Lemma 2.1.
There exist infinitely many positive integers $X$ such that
@@
where the set $\mathcal{A}$ is defined as in (2.1), depending on $X$.
To establish this result, we shall employ Harman’s sieve. Set
@@
Following usual terminology, we will refer to the sums in (2.7) as type I sums and the sums in (2.8) as type II sums. To establish Lemma 2.1, it now suffices to establish the bounds (2.7) and (2.8) above for $X\in\mathcal{X}$, where $\mathcal{X}$ is an infinite subset of $\mathbb{N}$. This set is constructed as follows. By the Dirichlet approximation theorem, mentioned in Section 1, there exist infinitely many positive integers $q$ such that
We shall take $\mathcal{X}$ to be the set of all positive integers $X$ such that
$X^{2\theta+10\eta}\leq q\leq X^{1-\theta-10\eta}$ (2.10)
for some $q\in\mathbb{N}$ satisfying (2.9), where $\eta$ is a suitably small positive number. After having established (2.7) and (2.8) for all $X\in\mathcal{X}$, Lemma 2.1 follows from Proposition 2.1. In the remainder of this article, we assume that $X\in\mathcal{X}$.
We conclude this section by deducing our main result, Theorem 1.2, from Lemma 2.1. Let $\gamma=1/c$ and $p\in\mathbb{O}$. Obviously, $p=\left[n^{c}\right]$ for some $n\in\mathbb{N}$ if and only if
Moreover, $(p+1)^{\gamma}=p^{\gamma}+\gamma p^{\gamma-1}+O\left(p^{\gamma-2}\right)$. Hence, if $\delta$ is defined as in (2.3), then
$p^{\gamma}+2\delta||<\delta.$ (2.13)
Therefore, Theorem 1.2 follows from Lemma 2.1.
@@ 3. Fourier analysis
We shall detect the conditions $\alpha a+\beta||<\Delta$ and $a^{\gamma}+2\delta||<\delta$ in (2.1) using the following Fourier analytic device.
@@
[…]
@@
To establish Theorem 1.2, it now suffices to prove that
@@ 4. Tailoring the type I and II sums
for $\eta>0$ small enough.
@@ 5. Estimations of $\mathcal{S}_{1}^{\ast}$ and $\mathcal{T}_{1}^{\ast}$
Recall from section 2 that $X\in\mathcal{X}$. Therefore, (2.9) and (2.10) hold for this $X$ and suitable $a,q\in\mathbb{Z}$. To estimate $\mathcal{S}_{1}^{\ast}$ and $\mathcal{T}_{1}^{\ast}$, defined in (4.1) and (4.4), we use the following standard bounds for sums involving linear exponential terms.
@@
[…]
@@
[…]
@@ 6. Estimations of $\mathcal{S}_{2}^{\ast}$ and $\mathcal{T}_{2}^{\ast}$
For the estimation of the type II sum $\mathcal{T}_{2}^{\ast}$, we will use the following bound due to Heath-Brown.
Lemma 6.2 (Heath-Brown).
Let $\mathcal{T}_{2}^{\ast}$ be given as in (4.5). Then, for arbitrary small real numbers $\varepsilon,\eta>0$, we have
$|\mathcal{T}_{2}^{\ast}|^{2}\ll$ $X^{2(\varepsilon+\eta)}\left(X^{5/2-\gamma}+X^{3-2\gamma}+X^{3-\gamma}N^{-1}+X^{(10-5\gamma)/3}N^{1/3}+X^{(8-4\gamma)/3}N^{2/3}+\right$.
$\left.X^{(10-8\gamma)/3}N^{4/3}\right).$
Proof.
The estimation of $\mathcal{T}_{2}^{\ast}$ is similar to the estimation of the term $L$ in [7, section 4]. The final estimate is found in [7, page 257] (with $X$ replaced by $N$, and NN replaced by $Y$, respectively). ∎
Our estimate of the type II sum $\mathcal{T}_{2}^{\ast}$ is as follows.
@@
Applying Lemma 6.2, we have
@@
where in the second line we have used that $MN\asymp X$. Thus $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$ if
@@
Recalling $X^{7/22}\leq M\leq X^{8/22}$ from (4.8) and $\gamma=1/c>8/9$ from Theorem 1.2, and noting that $1-\gamma<7/22<8/22<5\gamma-4$ if $\gamma>8/9$, the desired bound $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$ now follows upon taking $\varepsilon$ and $\eta$ sufficiently small. ∎
@@
Lemma 6.4 (Large $M$).
Suppose that $\varepsilon$ and $\eta$ are small enough and
$X^{5-5\gamma+6\varepsilon+18\eta}\ll M\ll X^{\gamma-2\varepsilon-6\eta}.$
Then $\mathcal{S}_{2}^{\ast}\ll X^{1-2\eta}$.
@@
provided that $M\leq X^{2/3}$. Now,
$X^{7/4-\gamma}M^{1/4}X^{\varepsilon+\eta}\ll X^{1-2\eta}\Longleftrightarrow M\ll X^{4\gamma-3-4\varepsilon-12\eta},$
$X^{2-\gamma}M^{-1/2}X^{\varepsilon+\eta}\ll X^{1-2\eta}\Longleftrightarrow M\gg X^{2-2\gamma+6\eta+2\varepsilon}$
and
$X^{15/8-\gamma}M^{-1/8}X^{\varepsilon+\eta}\ll X^{1-2\eta}\Longleftrightarrow M\gg X^{7-8\gamma+8\varepsilon+24\eta}.$
Since, $8/9<\gamma<1$ and $\varepsilon,\eta$ are arbitrary small, we have
$7-8\gamma+8\varepsilon+24\eta<2-2\gamma+6\eta+2\varepsilon.$
This implies the result in Lemma 6.5. ∎
@@
where we have used $V\leq H\ll X^{1-\gamma+\eta}$. This implies the result in Lemma $6.$ ∎
@@
if $\gamma>8/9$, Lemmas 6.4, 6.5 and 6.6 cover all relevant ranges if $\varepsilon$ and $\eta$ are small enough, and we thus have the following.
@@
Here we point out that the condition $\gamma>8/9$ above comes from the inequality $5-5\gamma<4\gamma-3$.
@@ 7. Estimations of $\mathcal{S}_{3}^{\ast}$ and $\mathcal{T}_{3}^{\ast}$
Lemma 7.2 (Large $M$).
Suppose that $\varepsilon$ and $\eta$ are small enough and
$X^{5-5\gamma+6\theta+8\varepsilon+20\eta}\ll M\ll X^{\gamma-2\theta-2\varepsilon-7\eta}.$
Then $\mathcal{S}_{3}^{\ast}\ll X^{1-2\eta}$.